# USTC-ICTS/PCFT-24-43

### v2

@@ IV.4.2 Approximate eigenvalues for $h\neq 2$
Recall that there are $d-1$ modes in the quadratic action (119) with $h\neq 2$. They turn out to have a much different 1-loop contribution. However, the difficulty of applying Sommerfeld-Watson resummation is that we do not have a simple expression for the eigen function and the eigenvalues. Nevertheless, for large enough eigenvalues, we may be able to find an approximate expression that captures the asymptotic behavior in the sum over infinite eigenvalues. Let us denote $\tilde{\mathcal{M}}_{\pm}$ as the set of approximate eigenvalues. The trick is that we can split the sum over eigenvalues into two parts. We can set up a subset $K_{\pm}\subset\mathcal{M}_{\pm}$ of eigenvalues. For eigenvalues $n\in K_{\pm}$ we use exact numeric solution, and for $n\notin K_{\pm}$ we use the approximation of large eigenvalues in $\tilde{\mathcal{M}}_{\pm}$ (see Fig. 7). For those eigenvalues in $\tilde{\mathcal{M}}_{\pm}$ not used in this approxiamtion, we denote them as $\tilde{K}_{\pm}$. In this treatment, the error is suppressed in the large size of $K_{\pm}$ limit.
@@
$\tilde{\mathcal{M}}_{+}$ =$=\{n\neq 0|\tan\frac{x}{2}\tan\frac{n}{4}+\frac{n}{\gamma x}=0\}$ (164)
$\tilde{\mathcal{M}}_{-}$ =$=\{n|\tan\frac{x}{2}\cot\frac{n}{4}-\frac{n}{\gamma x}=0\}$ (165)
where $n=0$ is excluded for $\tilde{\mathcal{M}}_{+}$ because it corresponds to a trivial $\tilde{f}_{+}$. On the other hand, for $\tilde{\mathcal{M}}_{-}$ we do not need to exclude a trivial eigenvalue like $n=2x$ as in $h=2$ case because there is no $n$ leading to trivial $\tilde{f}_{-}$. Therefore, $n=2x$ can regarded as an emergent mode for $h\rightarrow 2$. Even though we find two different types of $h$ for $d>9$ and $91<d<$ 9, the asymptotic eigenvalues $\tilde{\mathcal{M}}_{\pm}$ only depend on $\gamma$, which is smooth across $d=9$. Another feature of $\tilde{\mathcal{M}}_{\pm}$ is that the defining equation holds for $n\rightarrow-n$ just like the $h=2$ case. This is also the symmetry of full eigenfunction $f_{\pm}$, which means that we can double count the positive and negative eigenvalues in the Sommerfeld-Watson resummation with an additional $1/2$ factor as before.
@@
[…]
@@
Besides these nontrivial poles in (169) depending on $\tilde{K}_{\pm}$, the remaining poles are $z=0,\pm i\sqrt{\alpha}$ for $I_{+}$; and the remaining poles are $z=\pm i\sqrt{\alpha}$ for $I_{-}$. Using residue theorem at these points and integrate over $\alpha$, we find that  [^6] There is a subtlety when we have imaginary eigenvalues. For the imaginary eigenvalue $n_{*}$ in $\tilde{\mathcal{M}}_{+}$, we have $\sin^{2}(n_{*}/4)<0$, which implies $n_{*}^{2}+(\gamma x\tan\frac{x}{2})^{2}>0$. Therefore, when we do the integral over $\alpha$ in (169), it will hit a singularity of $1/F_{\pm}(\pm i\sqrt{\alpha})$ when ±$\pm i\sqrt{\alpha}=n_{*}$. To avoid this issue, note that the 1-loop determinant only cares about $|\sin^{2}(n/4)|$ and we can shift $n\rightarrow n+\epsilon$ in (168) for infinitesimal positive $\epsilon$ such that the average over $z=\pm n_{*}$ for $\log\frac{(z+\epsilon)^{2}}{(z+\epsilon)^{2}+(\gamma x\tan\frac{x}{2})^{2}}$ gives the real value $\log|\sin^{2}(n_{*}/4)|$. Adding this $\epsilon$ leads to the absolute value in the log term of (172). On the other hand, for the imaginary eigenvalue in $\tilde{\mathcal{M}}_{-}$, we have $\cos^{2}(n/4)>0$ and thus $n^{2}+(\gamma x\tan\frac{x}{2})^{2}<0$. It follows that the $\alpha$ integral does not hit any singularity and we do not need the $\epsilon$-prescription.
@@
IV.4.3 The first exact eigenvalue for $92\leq d<$ 9
Since we choose a nontrivial $\tilde{K}_{\pm}$, we need to include the exact eigenvalues in $K_{\pm}$ for compensation in (177). We will split the discussion into two parts: $h=1/2+i\nu$ ($92\leq d<$ 9) and $h<1/2$ ($d\geq 9$) because they have quite different behaviors. In this section, we will focus on the $h=1/2+i\nu$ case.
@@ IV.4.3 The first exact eigenvalue for $92\leq d<$ 9
Note that the RHS has unit norm, which leads to a real solution of $y$. Moreover, there are infinite numbers of solutions of $y$ because $y\rightarrow ye^{k\pi/\nu}$ for any $k\in\mathbb{Z}$ is also a solution. This means that as we decrease the temperature, there will be more and more pure imaginary eigenvalues. For our purpose, we find the critical temperature $1/\beta_{+}$ for the existence of the first pure imaginary eigenvalue corresponds to $k=-1$, namely
@@
where the power $(d-1)$ is due to $(d-1)$ numbers of $h\neq 2$ modes. As we restrict to $t>0$, here we should consider the validity of this partition function only for $\beta_{r}<\beta_{+}$. The inverse Laplace transformation for $E<0$  [^8] Here we momentarily take $E_{0}=0$ for simplicity. Recovering $E_{0}$ simply replaces $E\rightarrow E-E_{0}$. Note that the $\beta$ contour in the inverse Laplace transformation is to the left of $\beta_{+}\lambda^{-1/2}$ and leads to the exponential decay for $E<0$. leads to
@@
=$=\begin{cases}\frac{e^{\zeta}U(3/2-d/2,1-d/2,-\zeta)}{\Gamma((d-1)/2)}\quad&\zeta<0\\ \frac{U(-1/2,1-d/2,\zeta)}{\Gamma(3/2)}\quad&\zeta>0\end{cases}$ (201)
@@ V Thermodynmical meaning of the critical temperature
In previous sections, we have computed a critical temperature $T_{c}$ for the dcSYK model in the $q=0$ case and the $q\rightarrow 1$ case with $d<9$. As explained at the end of Sec. III.2.2, this critical temperature is read from the exponential tail of the spectrum and implies that we should regard the system as chaotic only for $T>T_{c}$. This physical interpretation has nothing to do with the critical temperature in the thermodynamical sense, defined as a non-smoothness change of free energy.
@@
The first is due to $-\log x$ being a convex function, and the second is the positivity of entropy. If the ensemble average gives a continuous spectrum in IR, the annealed free energy $F_{a}(1/T)$ will not be a monotonically decreasing function for small enough $T$. Therefore, the thermodynamically critical temperature is bounded from below
@@ VI.1 The spectrum
To check our analytic results, we compare them with results from exact diagonalization. In Fig.  12 we show the comparison between the numerical result of spectrum of the $N=15,~{}p=2$, $d=2,5$ model as well as the theoretical prediction of the double-scaled SYK model with parameter $\bar{q}$. We find good agreement between the bulk spectrum of analytic and numerical results. However, if we zoom in to the edge of the spectrum, we find a tail exceeding the theoretical prediction.
@@
We list the comparison between the theoretical value and the numerics in Table 2. We observe that for small $d$ the agreement is quite well, as $d$ increases the tail region becomes shorter so the fitting is less reliable. Examples of the fitting to the tail part is shown in Fig. 13.
@@
$T_{c}/\lambda^{1/2}$ $d=2$ $d=3$ $d=4$
Theory 0.1809 0.07717 0.03025
Numerics 0.1787 0.07667 0.03291
Error 1.2% 0.6% 8.8%
@@
$\tilde{E}_{0}$ $d=2$ $d=3$ $d=4$
Theory -3.4926 -3.0247 -2.8517
Numerics -3.5554 NA -3.3000
Error 1.8% NA 15.7%
@@ VI.2 The spectral form factor
In Fig. 14, we demonstrate the SFF for the dcSYK models with $N=12$ for $d=1$ and $d=2$, as well as the regular SYK model with random couplings normalized according to (7), which can be considered as the $d\to\infty$ limit of the dcSYK model. In each of the plots, the dashed lines are SFF for individual samples in the computation, and the solid curve is the average over all the samples. From the results, we observe that for $d=1$ the SFF does not contain a ramp and plateau even at infinite temperature $\beta=0$. On the contrary, in Fig. 14(b) for $d=2$ we observe the clear appearance of ramps and plateaus in the averaged SFF at $\beta=0$. As we decrease the temperature but still above the critical temperature $T_{c}$ (in the setup of Fig. 14, $T_{c}=0.187$, $\beta_{c}=5.35$ for $d=2$), the ramp becomes less steep comparing with the high temperature result. For example, as shown in Fig. 14(c) the slope of the ramp at $\beta=2$ is $k=0.45$ which is smaller than the slope of the ramp at $\beta=0$, which is $k=0.77$. As we decrease the temperature further below $T_{c}$ in Fig. 14(d), we observe big fluctuations in samples even at earlier time, this is qualitatively different from the regular SYK model as well as the dcSYK model at higher temperatures. Additionally, we do not observe increasing ramp in the SFF, in fact, if we fit the region from $10^{2}$ to $10^{3}$ in SFF, we observe a “down” ramp with a negative slope $k=-0.10$. In comparison, the SFF of the regular SYK model (6) at the same low temperature $\beta=10$ still exhibits clear ramp and plateau structure.
@@
(b) $d=2,\beta=0$, $k=0.77,h=2.4\times 10^{-4}$
(c) $d=2,\beta=2$, $k=0.45,h=4.5\times 10^{-4}$
(d) $d=2,\beta=10$, $h=2.1\times 10^{-2}$
(e) $d\to\infty$, $\beta=10$, $k=0.73,h=1.6\times 10^{-3}$
@@ VII Conclusion and discussions
We compute $T_{c}$ in two cases. The first case is $\lambda\rightarrow\infty$ where we can use free probability to find the exact form of the spectrum. The critical temperature $T_{c}=1/(2d)$ is read from the exponent of the exponential tail $e^{(E-E_{0})/T_{c}}$ for $E<E_{0}$ for some $E_{0}$. The exponential tail is from a non-pertrubative effect in $d$. The second case is $\lambda\rightarrow 0$ in the Schwarzian limit where we use the coarse-grained path integral to find the approximate spectrum in the IR. At the 1-loop level, we find the spectrum has an exponential tail for $d<9$. For $d>9$ we speculate the tail is due to higher loop effects. In both cases, we observe $T_{c}$ decreases as $d$ increases, which indicates a bigger chaotic regime. The thermodynamical meaning of the critical temperature $T_{c}$ is different for the two cases. For $\lambda\rightarrow\infty$, the thermodynamical critical temperature for a phase transition is much lower than $T_{c}$; for $\lambda\rightarrow 0$, it is of the same order as $T_{c}$.
@@
The existence of the critical temperature has a useful implication for future quantum simulation experiments for holographic systems (e.g. SYK model) under the slogan of “quantum gravity in a lab” initiated by Susskind in [62]. However, the regular SYK model has $O(N^{p})$ numbers of terms and is extremely hard to implement on a quantum simulator. On the other hand, the sparse SYK model [63] containing randomly picked $O(N)$ terms is much simpler and argued to have a holographic description as well. Along this line, one of the most interesting simulation projects is the teleportation protocol of traversable wormholes [64, 65, 66]. In a recent simulation [23], a small $N=7$ SYK model with only five terms in the Hamiltonian is learned to produce an arguable signal of traversable wormholes [67, 68] on Google’s Sycamore quantum chip. In this simulation, all terms in the learned Hamiltonian are commutative to each other and the model is integrable though it is slightly different from our construction (3) for $d=1$. The analysis in [24] shows that for small $N$ one cannot clearly distinguish the simulation signals from the traversable wormhole mechanism in the chaotic regular SYK model or the thermalization process in the integrable commuting SYK model (see also [69]).

### v4

@@ IV.4.2 Approximate eigenvalues for $h\neq 2$
Recall that there are d1d-1 modes in the quadratic action (119) with $h\neq 2$. They turn out to have a much different 1-loop contribution. However, the difficulty of applying Sommerfeld-Watson resummation is that we do not have a simple expression for the eigen function and the eigenvalues. Nevertheless, for large enough eigenvalues, we may be able to find an approximate expression that captures the asymptotic behavior in the sum over infinite eigenvalues. Let us denote $\tilde{\mathcal{M}}_{\pm}$ as the set of approximate eigenvalues. The trick is that we can split the sum over eigenvalues into two parts. We can set up a subset $K_{\pm}\subset\mathcal{M}_{\pm}$ of eigenvalues. For eigenvalues $n\in K_{\pm}$ we use exact numeric solution, and for $n\notin K_{\pm}$ we use the approximation of large eigenvalues in $\tilde{\mathcal{M}}_{\pm}$ (see Fig. 7). For those eigenvalues in $\tilde{\mathcal{M}}_{\pm}$ not used in this approxiamtion, we denote them as $\tilde{K}_{\pm}$. In this treatment, the error is suppressed in the large size of $K_{\pm}$ limit.
@@
$\tilde{\mathcal{M}}_{+}$ $=\{n\neq 0|\tan\frac{x}{2}\tan\frac{n}{4}+\frac{n}{\gamma x}=0\}$ (164)
$\tilde{\mathcal{M}}_{-}$ $=\{n|\tan\frac{x}{2}\cot\frac{n}{4}-\frac{n}{\gamma x}=0\}$ (165)
where $n=0$ is excluded for $\tilde{\mathcal{M}}_{+}$ because it corresponds to a trivial $\tilde{f}_{+}$. On the other hand, for $\tilde{\mathcal{M}}_{-}$ we do not need to exclude a trivial eigenvalue like $n=2x$ as in $h=2$ case because there is no $n$ leading to trivial $\tilde{f}_{-}$. Therefore, $n=2x$ can regarded as an emergent mode for $h\rightarrow 2$. Even though we find two different types of $h$ for $d>9$ and $1<d<9$, the asymptotic eigenvalues $\tilde{\mathcal{M}}_{\pm}$ only depend on $\gamma$, which is smooth across $d=9$. Another feature of $\tilde{\mathcal{M}}_{\pm}$ is that the defining equation holds for $n\rightarrow-n$ just like the $h=2$ case. This is also the symmetry of full eigenfunction $f_{\pm}$, which means that we can double count the positive and negative eigenvalues in the Sommerfeld-Watson resummation with an additional $1/2$ factor as before.
@@
$I_{\pm}=-\frac{1}{4\pi i}\int_{0}^{\gamma^{2}x^{2}\tan^{2}x/2}d\alpha\int_{\mathcal{C}_{\pm}}dz\frac{1}{\alpha+z^{2}}\frac{F_{\pm}^{\prime}(z)}{F_{\pm}(z)}$ (169)
@@
Besides these nontrivial poles in (169) depending on $\tilde{K}_{\pm}$, the remaining poles are $z=0,\pm i\sqrt{\alpha}$ for $I_{+}$; and the remaining poles are $z=\pm i\sqrt{\alpha}$ for $I_{-}$. Using residue theorem at these points and integrate over $\alpha$, we find that  [^6] There is a subtlety when we have imaginary eigenvalues. For the imaginary eigenvalue $n_{*}$ in $\tilde{\mathcal{M}}_{+}$, we have $\sin^{2}(n_{*}/4)<0$, which implies $n_{*}^{2}+(\gamma x\tan\frac{x}{2})^{2}>0$. Therefore, when we do the integral over $\alpha$ in (169), it will hit a singularity of $1/F_{\pm}(\pm i\sqrt{\alpha})$ when $\pm i\sqrt{\alpha}=n_{*}$. To avoid this issue, note that the 1-loop determinant only cares about $|\sin^{2}(n/4)|$ and we can shift $n\rightarrow n+\epsilon$ in (168) for infinitesimal positive $\epsilon$ such that the average over $z=\pm n_{*}$ for $\log\frac{(z+\epsilon)^{2}}{(z+\epsilon)^{2}+(\gamma x\tan\frac{x}{2})^{2}}$ gives the real value $\log|\sin^{2}(n_{*}/4)|$. Adding this $\epsilon$ leads to the absolute value in the log term of (172). On the other hand, for the imaginary eigenvalue in $\tilde{\mathcal{M}}_{-}$, we have $\cos^{2}(n/4)>0$ and thus $n^{2}+(\gamma x\tan\frac{x}{2})^{2}<0$. It follows that the $\alpha$ integral does not hit any singularity and we do not need the $\epsilon$-prescription.
@@
IV.4.3 The first exact eigenvalue for $2\leq d<9$
Since we choose a nontrivial $\tilde{K}_{\pm}$, we need to include the exact eigenvalues in $K_{\pm}$ for compensation in (177). We will split the discussion into two parts: $h=1/2+i\nu$ ($2\leq d<9$) and $h<1/2$ ($d\geq 9$) because they have quite different behaviors. In this section, we will focus on the $h=1/2+i\nu$ case.
@@ IV.4.3 The first exact eigenvalue for $2\leq d<9$
Note that the RHS has unit norm, which leads to a real solution of $y$. Moreover, there are infinite numbers of solutions of $y$ because $y\rightarrow ye^{k\pi/\nu}$ for any $k\in\mathbb{Z}$ is also a solution. This means that as we decrease the temperature, there will be more and more pure imaginary eigenvalues. For our purpose, we find the critical temperature $1/\beta_{+}$ for the existence of the first pure imaginary eigenvalue corresponds to k=1k=-1, namely
@@
where the power (d1)(d-1) is due to (d1)(d-1) numbers of $h\neq 2$ modes. As we restrict to $t>0$, here we should consider the validity of this partition function only for $\beta_{r}<\beta_{+}$. The inverse Laplace transformation for $E<0$  [^8] Here we momentarily take $E_{0}=0$ for simplicity. Recovering $E_{0}$ simply replaces $E\rightarrow E-E_{0}$. Note that the $\beta$ contour in the inverse Laplace transformation is to the left of $\beta_{+}\lambda^{-1/2}$ and leads to the exponential decay for $E<0$. leads to
@@
$=\begin{cases}\frac{e^{\zeta}U(3/2-d/2,1-d/2,-\zeta)}{\Gamma((d-1)/2)}\quad&\zeta<0\\ \frac{U(-1/2,1-d/2,\zeta)}{\Gamma(3/2)}\quad&\zeta>0\end{cases}$ (201)
@@ V Thermodynmical meaning of the critical temperature
In previous sections, we have computed a critical temperature $T_{c}$ for the dcSYK model in the $q=0$ case and the $q\rightarrow 1$ case with $d<9$. Since this critical temperature is read from the exponent of the infinite exponential tail of the spectrum, we should naturally expect that there does not exist quantum chaos for $T\lessapprox T_{c}$ because the system simply does not have enough eigenvalues to scramble in the tail regime with a sparse spectrum. On the other hand, the existence of the $h=2$ mode for $q\rightarrow 1$, which is identical to the Schwarzian mode, suggests that quantum chaos should exist for $T\gtrapprox T_{c}$. This transition of quantum chaos will be further supported by our numerical studies of spectral form factor in Sec. VI.2. However, this physical interpretation has nothing to do with the critical temperature in the thermodynamical sense, defined as a non-smoothness change of free energy.
@@
The first is due to $\log x$ being a convex function, and the second is the positivity of entropy. If the ensemble average gives a continuous spectrum in IR, the annealed free energy $F_{a}(1/T)$ will not be a monotonically decreasing function for small enough $T$. Therefore, the thermodynamically critical temperature is bounded from below
@@ VI.1 The spectrum
To check our analytic results, we compare them with results from exact diagonalization. In Fig.  12 we show the comparison between the numerical result of spectrum of the N=15,p=2N=15,~p=2, $d=2,5$ model as well as the theoretical prediction of the double-scaled SYK model with parameter $\bar{q}$. We find good agreement between the bulk spectrum of analytic and numerical results. However, if we zoom in to the edge of the spectrum, we find a tail exceeding the theoretical prediction.
@@
We can compare this with the numerical results. Examples of the fitting to the tail part is shown in Fig. 13. We also list the comparison between the theoretical value and the numerics in Table 2. We observe that for small $d$ the agreement is quite well, as $d$ increases the tail region becomes shorter so the fitting is less reliable.
@@
$T_{c}/\lambda^{1/2}$ $d=2$ $d=3$
Theory 0.1809 0.07717
Numerics 0.1836 0.07709
Error 1.5% 0.1%
@@
$\tilde{E}_{0}$ $d=2$ $d=3$
Theory -3.4926 -3.0247
Numerics -3.5606 NA
Error 1.9% NA
@@ VI.2 The spectral form factor
In Fig. 14, we demonstrate the SFF for the dcSYK models with $N=12$ for $d=1$ and $d=2$, as well as the regular SYK model with random couplings normalized according to (7), which can be considered as the $d\to\infty$ limit of the dcSYK model. In each of the plots, the dashed lines are SFF for individual samples in the computation, and the solid curve is the average over all the samples. From the results, we observe that for $d=1$ the SFF does not contain a ramp and plateau even at infinite temperature $\beta=0$. On the contrary, in Fig. 14(b) for $d=2$ we observe the clear appearance of ramps and plateaus in the averaged SFF at $\beta=0$. As we decrease the temperature but still above the critical temperature $T_{c}$ (in the setup of Fig. 14, $T_{c}=0.187$, $\beta_{c}=5.35$ for $d=2$), the ramp becomes less steep comparing with the high temperature result. For example, as shown in Fig. 14(c) the slope of the ramp at $\beta=2$ is $k=0.45$ which is smaller than the slope of the ramp at $\beta=0$, which is $k=0.77$. As we decrease the temperature further below $T_{c}$ in Fig. 14(d), we observe big fluctuations in samples even at earlier time, this is qualitatively different from the regular SYK model as well as the dcSYK model at higher temperatures. Additionally, we do not observe increasing ramp in the SFF, in fact, if we fit the region from $10^{2}$ to $10^{3}$ in SFF, we observe a “down” ramp with a negative slope k=0.10k=-0.10. In comparison, the SFF of the regular SYK model (6) at the same low temperature $\beta=10$ still exhibits clear ramp and plateau structure.
@@
(b) $d=2,\beta=0$, $k=0.74,h=2.4\times 10^{-4}$
(c) $d=2,\beta=2$, $k=0.45,h=4.3\times 10^{-4}$
(d) $d=2,\beta=10$, $h=2.0\times 10^{-2}$
(e) $d\to\infty$, $\beta=10$, $k=0.84,h=1.6\times 10^{-3}$
@@ VII Conclusion and discussions
We compute $T_{c}$ in two cases. The first case is $\lambda\rightarrow\infty$ where we can use free probability to find the exact form of the spectrum. The critical temperature $T_{c}=1/(2d)$ is read from the exponent of the exponential tail $e^{(E-E_{0})/T_{c}}$ for $E<E_{0}$ for some $E_{0}$. The exponential tail is from a non-perturbative effect in $d$. The second case is $\lambda\rightarrow 0$ in the Schwarzian limit where we use the coarse-grained path integral to find the approximate spectrum in the IR. At the 1-loop level, we find the spectrum has an exponential tail for $d<9$. For $d>9$ we speculate the tail is due to higher loop effects. In both cases, we observe $T_{c}$ decreases as $d$ increases, which indicates a bigger chaotic regime. The thermodynamical meaning of the critical temperature $T_{c}$ is different for the two cases. For $\lambda\rightarrow\infty$, the thermodynamical critical temperature for a phase transition is much lower than $T_{c}$; for $\lambda\rightarrow 0$, it is of the same order as $T_{c}$.
@@
The existence of the critical temperature has a useful implication for future quantum simulation experiments for holographic systems (e.g. SYK model) under the slogan of “quantum gravity in a lab” initiated by Susskind in [62]. However, the regular SYK model has $O(N^{p})$ numbers of terms and is extremely hard to implement on a quantum simulator. On the other hand, the sparse SYK model [63] containing randomly picked $O(N)$ terms is much simpler and argued to have a holographic description as well. Along this line, one of the most interesting simulation projects is the teleportation protocol of traversable wormholes [64, 65, 66]. In a recent simulation [23], an attempt was made with a small $N=7$ SYK model with only five terms in the Hamiltonian that is learned to produce an arguable signal of traversable wormholes [67, 68] on Google’s Sycamore quantum chip. In this simulation, all terms in the learned Hamiltonian are commutative to each other and the model is integrable though it is slightly different from our construction (3) for $d=1$. The analysis in [24] shows that for small $N$ one cannot clearly distinguish the simulation signals from the traversable wormhole mechanism in the chaotic regular SYK model or the thermalization process in the integrable commuting SYK model (see also [69]).
@@
  4. 4.
    Another diagonose of quantum chaos is the out-of-time-order correlators (OTOC). Since our model demonstrates a chaotic-integrable transition, it is interesting to check if we can observe similar transition in OTOC. On the other hand, since there are examples of OTOC overestimating the chaotic behavior of special models with unstable fixed points [75, 76, 77, 78, 79], it is also interesting to test this in our model.